# https://vision.ist.i.kyoto-u.ac.jp/

### v2

@@ Abstract
In this paper, we introduce a novel training method for making any monocular depth network learn absolute scale and estimate metric road-scene depth just from regular training data, i.e., driving videos. We refer to this training framework as StableCamH. The key idea is to leverage cars found on the road as sources of scale supervision but to incorporate them in the training robustly. StableCamH detects and estimates the sizes of cars in the frame and aggregates scale information extracted from them into a camera height estimate whose consistency across the entire video sequence is enforced as scale supervision. This realizes robust unsupervised training of any, otherwise scale-oblivious, monocular depth network to become not only scale-aware but also metric-accurate without the need for auxiliary sensors and extra supervision. Extensive experiments on the KITTI and Cityscapes datasets show the effectiveness of StableCamH and its state-of-the-art accuracy compared with related methods. We also show that StableCamH enables training on mixed datasets of different camera heights, which leads to larger-scale training and thus higher generalization. Metric depth reconstruction is essential in any road-scene visual modeling, and StableCamH democratizes its deployment by establishing the means to train any model as a metric depth estimator.
@@ 1 Introduction
Monocular depth estimation is essential for autonomous driving and advanced driver assistance systems (ADAS). It underlies many of their key perceptual tasks including 3D object detection [7, 41, 52, 34], motion planning [40, 54, 33], and road segmentation [28]. Although supervised methods have achieved impressive accuracy, they suffer from costly ground-truth data collection. This hinders their generalizability [35] and also makes continuous learning difficult.
@@
Many recent methods [5, 55, 27, 15, 18] leverage self-supervision to avoid the need for costly supervision. Most of these methods, however, suffer from scale ambiguity which renders them difficult to use off-the-shelf. To resolve the scale ambiguity and recover the metric scale, the actual scene scale needs to be supervised in some form. Past works have adopted velocity [16], GPS [3], IMU [56, 43], or camera height [38, 48] for this. These methods require such auxiliary sensors in addition to the monocular RGB camera which significantly limits the training data. Arbitrary in-the-wild driving videos cannot be used as they lack such measurements. These self-supervised methods still require weak scale supervision and cannot be made fully self-supervised as the scale changes for each sequence captured with a different camera.
@@
Roads, however, have plenty of objects whose projections in the video frames can inform the metric scale. Most conveniently, cars are common objects on the road on which we would like to reconstruct the depth accurately in metric scale the most. If we can somehow leverage the known sizes of cars just from their projections in the video frames, we may train a mono-depth network to be metric accurate from arbitrary driving videos. Simply learning from an object size prior would, however, be too brittle since the metric supervision will be as ambiguous as the accuracy of that prior, which likely cannot be made perfect. Not only the perspective projection, but also the variations in car makes, years, and colors, cause large variance in the size estimates, which will be directly reflected in the mono-depth errors. In addition, there are no reliable object size priors except for a few introduced as probability densities for other tasks [60, 20, 57], whose parameters are challenging to find.
@@
How can we leverage prior knowledge about object sizes in a way robust to the actual instantiations in the scene? We derive a novel training framework which we refer to as StableCamH that answers this question. Our key idea is to aggregate object size cues in the frames and sequence into a single physical measure, the camera height. Regardless of the scene, the camera capturing it is the same for the same sequence. This means that if we can transform the object size cues in different frames into a camera height estimate, the very fact that it should not change can be used as a powerful invariance supervision to achieve metric depth reconstruction. We can formulate this as an optimization of the camera height across frames and training epochs. This leads to gradual metric-aware learning that is stable and accurate, which also makes the depth estimates consistent across the sequence.
@@
For estimating the car sizes in each frame, we introduce Learned Size Prior (LSP). This prior estimates vehicle dimensions from its appearance and is trained with various augmentations on a large-scale dataset collected by web scraping without manual annotations. By comparing these dimension estimates in a frame with the ones computed from the estimated depth map, we obtain a per-frame scale factor. This scale can be multiplied with each camera height derived from the depth map to obtain the metric camera height for each frame. These estimated camera heights across all frames in a sequence are then consolidated by their median and used to update the supervision with weighted moving average at every end of training epochs.
@@
We evaluate the effectiveness of StableCamH with extensive experiments using the KITTI and Cityscapes datasets and compare its accuracy with a number of weakly-supervised methods. The results clearly show that StableCamH succeeds in making monocular depth estimation methods learn metric scales and achieves the state-of-the-art accuracy despite the lack of depth and scale supervision. We also demonstrate its capability to train a model on mixed datasets collected with different cameras, i.e., camera heights. This enables larger scale training and leads to higher generalization. We believe StableCamH can serve as a fundamental building block for monocular depth estimation in the wild.
@@ 2.1 Self-supervision
Although, as in other tasks, supervised methods achieve the highest accuracy in monocular depth estimation (MDE) [1, 29, 24, 53, 31], they obviously require ground-truth data usually obtained with a LiDAR sensor. Self-supervised MDE methods remove the need for such hardware costs for training. Garg et al. [11] proposed the first approach to self-supervised MDE by predicting per-pixel disparities between a stereo pair. Godard et al. [14] extend this idea by introducing left-right consistency and greatly improve the accuracy. These methods require calibrated stereo data for training which limits their applicability. Zhou et al. [59] introduced a pioneering method that enables training solely from monocular videos by incorporating a network that estimates the related camera pose between successive frames. Based on this idea, many MDE methods trainable with monocular videos have been proposed [15, 27, 58, 55, 5, 18].
Despite the popularity and improvements of self-supervised MDE, directly deploying those models in autonomous driving or ADAS systems is unrealistic since they can only produce relative depth, up to an unknown scale. Our StableCamH overcomes this issue without the need for any external supervision.
@@ 2.2 Weak Supervision
Several works have tackled scale-aware MDE by adopting weak scale supervision through velocity [16], GPS [3], IMU [43], or camera height [38, 48] measurements. Zhang et al. [56] use the combination of IMU, velocity, and gravity to achieve the highest accuracy among these methods. These methods fundamentally rely on such auxiliary measurements in addition to regular RGB images. This makes it impossible to take advantage of the many road-scene videos on the Internet (e.g., YouTube) and easily installed dash cam for training. In contrast, StableCamH relies on a universal object size prior which is independent of the data collection environments and does not require any such weak supervision; it can be trained solely from monocular video sequences.
@@ 2.3 Object Size Priors
Several works for other tasks leverage priors on object sizes including single view metrology [60], object insertion [20, 57], and visual odometry [10, 36]. These methods model the priors as normal distributions or Gaussian mixture models, and the parameters of these distributions are predefined. When there is a gap between those values and the true ones, which is usually the case in the real world, the difference is directly reflected in erroneous scale estimates.
Instead of these fixed priors, we introduce a learning-based object size prior dubbed LSP. This estimates vehicle dimensions based on its appearance, thus can avoid the aforementioned problem. Furthermore, it is trained on a large-scale dataset with various augmentations, so that it can robustly estimate vehicle dimensions from any image.
@@ 3 Preliminaries
Our goal is to train a metrically accurate MDE model $\theta_{0}pt$ with only self-supervision. This self-supervision can be subdivided into two parts: one designed for learning metric scale (Sec. 4) and the other dedicated to learning depth geometry. For the latter, we follow Zhou et al. [59] and minimize the reconstruction loss between the current frame $I_{t}$ and the ones synthesized from temporally adjacent frames $I_{s},(s=t\pm 1)$. We employ a pose network $\theta_{\mathrm{pose}}$ in addition to $\theta_{0}pt$ to estimate the relative camera pose $\mathbf{T}_{t\rightarrow s}\in\mathrm{SE}(3)$ and the current dense depth map $D_{t}$, respectively:
@@
We then synthesize the image from the current viewpoint $I_{s\rightarrow t}$:
@@
where $\expectationvalue{\cdot}$ is a bilinear sampling operator, $proj(\cdot)$ outputs the 2D coordinates of $D_{t}$ when projected into the viewpoint of $I_{s}$, and $\mathbf{K}$ is the camera intrinsics. Following Godard et al. [15], we use an affine combination of $\mathrm{SSIM}$ [44] and L1 loss as the photometric error $pe$ and define the reconstruction loss $\mathcal{L}_{\mathrm{rec}}$ as the photometric error between $I_{t}$ and $I_{s\rightarrow t}$ for the pixels with minimum error across source frames:
@@
where $d^{*}_{t}=d_{t}\big{/}\bar{d_{t}}$ is the mean-normalized inverse depth [39] to prevent the estimated depth from shrinking. Furthermore, when computing $\mathcal{L}_{\mathrm{rec}}$ we use auto-masking [15] to mask out stationary pixels, which generate incorrect supervisory signals.
Figure 1: Overview of StableCamH. At each training step $n$, an unscaled camera height $H^{\prime n,\tau}_{\mathrm{cam}}$ is differentiablly computed from the estimated depth. The previous epoch $\tau-1$ provides supervision with a scaled camera height $H^{*\tau-1}_{\mathrm{cam}}$. To compute this scaled camera height supervision, Silhouette Projector first computes the object silhouette heights $H^{\prime}_{\mathrm{obj}}$ from the depth map. By comparing $H^{\prime}_{\mathrm{obj}}$ and the estimated one $H_{\mathrm{obj}}$ with LSP, per-frame scale factor $s$ is determined and we obtain the scaled camera height $H^{n,\tau}_{\mathrm{cam}}=s\cdot H^{\prime n,\tau}_{\mathrm{cam}}$. At the end of each epoch, $H^{*}_{\mathrm{cam}}$ is optimized across a series of consecutive frames and updated with weighted moving average.
4 StableCamH
Fig. 1 depicts the overall architecture of StableCamH. Please see the caption for an overview of the framework. Let “$\,{}^{\prime}\,$” denote “not metrically scaled”.
@@ 4.1 Scale-Aware Self-Supervised Learning
In StableCamH, we derive a camera height pseudo-supervision from depth estimates. To obtain the per-frame camera height, we first compute the normal vector for each pixel $p_{i}$. We consider the 8-neighborhood of $p_{i}$ and form 8 pairs $\mathcal{N}(p_{i})=\{(p^{i}_{j_{0}},p^{i}_{j_{1}})\}^{8}_{j=1}$. Each pair $(p^{i}_{j_{0}},p^{i}_{j_{1}})$ consists of two pixels whose difference vectors w.r.t. $p_{i}$ are perpendicular to each other, and are sorted counterclockwise. We reproject the paired pixels onto 3D space and compute the cross-product of the two difference vectors w.r.t. the reprojected 3D point of $p_{i}$. By summing and normalizing all the cross-product vectors of each pair, we obtain the normal vector $\hat{\mathbf{n}}(p_{i})$
@@
where $p_{i}$ is in the road region mask $M_{\mathrm{r}}$ obtained with an off-the-shelf semantic segmentation model [23] before training. In our training framework, the networks learn the real-world scale with the pseudo camera height, so we need the camera height in absolute scale. The camera height $\tilde{H}^{\prime}_{\mathrm{cam}}$ computed from the estimated depth map is, however, unscaled due to the lack of scale constraints from the reconstruction loss. By using Silhouette Projector, we can aggregate scale information from the object height prior into a per-frame scale factor $s$ (see Sec. 4.2) and obtain the scaled camera height $H_{\mathrm{cam}}=s\cdot\tilde{H}^{\prime}_{\mathrm{cam}}$.
The scaled camera heights $H_{\mathrm{cam}}$ computed for individual frames are inherently inaccurate due to imperfections in the estimated depth maps and the size priors. We leverage the fact that the camera height does not change in a sequence and formulate the invariance into an optimization across frames and training epochs. Thanks to this optimization, we acquire an accurate and stable supervision, which is also consistent across the scene. At the end of the $\tau$-th epoch, we use the median of $\{H^{n,\tau}_{\mathrm{cam}}\}^{N}_{n=1}$ for all frames $N$ as the representative camera height $H^{\tau}_{\mathrm{cam}}$ at the $\tau$-th epoch. Then, we update the scaled camera height pseudo-supervision $H^{*}_{\mathrm{cam}}$ with its weighted moving average:
@@
When training on a dataset containing images captured at multiple camera heights, we create pseudo-labels corresponding to each camera height and individually optimize them.
To make the network learn the absolute scale and directly produce the scaled depth map, we impose a loss on the camera height $\mathcal{L}_{\mathrm{cam}}$ and supervise the camera height $H^{\prime}_{\mathrm{cam}}$ computed from the estimated depth map with $H^{*\tau-1}_{\mathrm{cam}}$:
@@
Note that we use $H^{*\tau-1}_{\mathrm{cam}}$ as the pseudo-supervision during the $\tau$-th epoch training since $H^{*\tau}_{\mathrm{cam}}$ is computed at the end of this epoch.
It is worth noting that at inference we do not require either the camera height invariance assumption in a sequence or its optimization, since we employ these in order to obtain the accurate pseudo-supervision during training. The computational cost at inference is exactly the same as an original MDE model trained without StableCamH.

### v3

@@ Abstract
In this paper, we introduce a novel training method for making any monocular depth network learn absolute scale and estimate metric road-scene depth just from regular training data, i.e., driving videos. We refer to this training framework as FUMET. The key idea is to leverage cars found on the road as sources of scale supervision and to incorporate them in network training robustly. FUMET detects and estimates the sizes of cars in a frame and aggregates scale information extracted from them into an estimate of the camera height whose consistency across the entire video sequence is enforced as scale supervision. This realizes robust unsupervised training of any, otherwise scale-oblivious, monocular depth network so that they become not only scale-aware but also metric-accurate without the need for auxiliary sensors and extra supervision. Extensive experiments on the KITTI and the Cityscapes datasets show the effectiveness of FUMET, which achieves state-of-the-art accuracy. We also show that FUMET enables training on mixed datasets of different camera heights, which leads to larger-scale training and better generalization. Metric depth reconstruction is essential in any road-scene visual modeling, and FUMET democratizes its deployment by establishing the means to convert any model into a metric depth estimator.
@@ 1 Introduction
Monocular depth estimation is essential for autonomous driving and advanced driver assistance systems (ADAS). It underlies many of their key perceptual tasks including 3D object detection [10, 45, 56, 38], motion planning [44, 58, 36], and road segmentation [31]. Although supervised methods have achieved impressive accuracy, they suffer from costly ground-truth data collection. This limits their generalizability [39] and also makes continuous learning difficult.
@@
Many recent methods [8, 59, 30, 18, 21] leverage self-supervision to scale up training. Most of these methods, however, suffer from scale ambiguity which renders them difficult to use off-the-shelf. To resolve the scale ambiguity and recover in metric scale, the actual scene scale needs to be supervised in some form. Past works have leveraged velocity [19], GPS [6], IMU [60, 47], or camera height [42, 52] for this. These methods require auxiliary sensors to measure these extra quantities in addition to the monocular RGB camera, which significantly limits the ability to scale the size of training data. Arbitrary in-the-wild driving videos cannot be used as they lack such measurements. These self-supervised methods often still require weak scale supervision and cannot be made fully self-supervised as the scale changes for each sequence captured with a different camera.
@@
Road scenes actually have plenty of untapped sources of metric scale. Cars found on the road are one of them as they are rigid objects whose actual sizes do not change and are unique to each make and model. The road is also exactly the region where we need accurate metric depth the most. If we can leverage the sizes of cars to bridge the 3D to 2D projection, we may train a mono-depth network to become metric-accurate. Simply employing a prior on the car size, however, would be too brittle since the metric supervision will be as ambiguous as the accuracy of that prior. The perspective projection and the variations in car makes, years, and colors, lead to a large variance in the size estimates, which will directly affect the mono-depth metric accuracy. In addition, there are no reliable object size priors except for a few introduced as probability densities for other tasks [64, 23, 61] whose parameters are not shared.
@@
How can we train an arbitrary mono-depth estimator to become metric-accurate by robustly leveraging prior knowledge about the 3D sizes of cars found in driving videos? We derive fully unsupervised metric depth trainer (FUMET) to answer this question. Our key idea is to aggregate car size cues in the frames and the sequence into a single physical measure, namely the camera height. Regardless of the scene, the camera capturing it is the same for the same sequence. This means that, by transforming the car size cues found in different frames into a camera height estimate, the very fact that it should not change can be used as a supervision of invariance to achieve metric depth reconstruction. We can formulate this as an optimization of the camera height across frames and training epochs. This leads to gradual metric-aware learning that is stable and accurate, which also makes the depth estimates consistent across the sequence.
@@
We introduce a learned size prior (LSP) to estimate the size of each car found in a frame. This prior gives us the vehicle dimensions from its appearance. It is trained with various augmentations on a large-scale dataset without the need for any manual annotation. By comparing these dimension estimates with those computed from the depth estimates, we obtain a per-frame scale factor. The multiplication of this scale estimate with the camera height estimate derived from the estimated depth becomes the camera height estimate of the frame. These estimated camera heights across all frames in a sequence are then consolidated by using their median value which is then used as scale supervision with its weighted moving average at the end of each training epoch.
@@
We evaluate the effectiveness of FUMET with extensive experiments using the KITTI and the Cityscapes datasets and compare its accuracy with a number of weakly-supervised methods. The results clearly show that FUMET succeeds in making monocular depth estimation methods learn metric scales and achieves the state-of-the-art accuracy without any direct depth and scale supervision. We also demonstrate its capability to train a model on mixed datasets collected with different cameras, i.e., camera heights. This enables larger-scale training and leads to higher generalization. We believe FUMET can serve as a fundamental building block for monocular depth estimation in the wild.
@@ 2.1 Self-Supervision
Although, as in other tasks, supervised methods achieve the highest accuracy in monocular depth estimation (MDE) [3, 32, 27, 57, 34], they obviously require ground-truth data usually obtained with a LiDAR sensor. Self-supervised MDE methods remove the need for such hardware costs for training. Garg et al. [14] proposed the first approach to self-supervised MDE by predicting per-pixel disparities between a stereo pair. Godard et al. [17] extend this idea by introducing left-right consistency and greatly improve the accuracy. These methods require calibrated stereo data for training which limits their applicability. Zhou et al. [63] introduced a pioneering method for training solely from monocular videos by incorporating a network that estimates the relative camera poses between successive frames. Based on this idea, many MDE methods that can be trained solely with monocular videos have been proposed [18, 30, 62, 59, 8, 21].
Despite the popularity and improvements of self-supervised MDE, directly deploying those models in autonomous driving or ADAS systems is unrealistic as they only produce relative depth up to an unknown scale. Our FUMET overcomes this issue without the need for any external supervision.
@@ 2.2 Weak Supervision
Several works have tackled scale-aware MDE by adopting weak scale supervision through velocity [19], GPS [6], IMU [47], or camera height [42, 52] measurements. Zhang et al. [60] use the combination of IMU, velocity, and gravity to achieve the highest accuracy among these methods. These methods fundamentally rely on such auxiliary measurements in addition to regular RGB images. This makes it impossible to take advantage of the many road-scene videos on the Internet (e.g., YouTube) or casually installed dashcam videos for training. In contrast, FUMET relies on a universal car size prior which is independent of the data collection environments and does not require any such weak supervision; it can be trained solely from monocular video sequences.
@@ 2.3 Object Size Priors
Several works for other tasks leverage priors on object sizes including single view metrology [64], object insertion [23, 61], and visual odometry [13, 40]. These methods model the priors as normal distributions or Gaussian mixture models, and the parameters of these distributions are predetermined. When there is a gap between those values and the true ones, which is usually the case in the real world, the difference directly surfaces in erroneous scale estimates.
Instead of these fixed priors, we introduce a learning-based size prior which we refer to as LSP. This estimates vehicle dimensions based on its appearance, thus can avoid the aforementioned problem. Furthermore, it is trained on a large-scale dataset with various augmentations, so that it can robustly estimate vehicle dimensions from any image.
@@ 3 Preliminaries
Our goal is to train a metric-accurate MDE model $\theta_{0}pt$ with only self-super-vision. This self-supervision is comprised of two parts: one designed for learning metric scale (Sec. 4) and the other dedicated to learning depth geometry. For the latter, we follow Zhou et al. [63] and minimize the reconstruction loss between the current frame $I_{t}$ and the ones synthesized from temporally adjacent frames $I_{s},(s=t\pm 1)$. We employ a pose network $\theta_{\mathrm{pose}}$ in addition to $\theta_{0}pt$ to estimate the relative camera pose $\mathbf{T}_{t\rightarrow s}\in\mathrm{SE}(3)$ and the current dense depth map $D_{t}$, respectively,
@@
We then synthesize the image from the current viewpoint
@@
where $\expectationvalue{\cdot}$ is a bilinear sampling operator, $proj(\cdot)$ outputs the 2D coordinates of $D_{t}$ when projected into the viewpoint of $I_{s}$, and $\mathbf{K}$ is the camera intrinsics. Following Godard et al. [18], we use an affine combination of $\mathrm{SSIM}$ [48] and L1 loss as the photometric error $pe$ and define the reconstruction loss $\mathcal{L}_{\mathrm{rec}}$ as the photometric error between $I_{t}$ and $I_{s\rightarrow t}$ for the pixels with minimum error across source frames
@@
where $d^{*}_{t}=d_{t}\big{/}\bar{d_{t}}$ is the mean-normalized inverse depth [43] to prevent the estimated depth from shrinking. Furthermore, when computing $\mathcal{L}_{\mathrm{rec}}$ we use auto-masking [18] to mask out stationary pixels which generate incorrect supervisory signals.
Figure 1: Overview of FUMET. At each training step $n$, an unscaled camera height $H^{\prime n,\tau}_{\mathrm{cam}}$ is computed differentiablly from the estimated depth. The previous epoch $\tau-1$ provides supervision with a scaled camera height $H^{*\tau-1}_{\mathrm{cam}}$. To obtain this scaled camera height supervision, Silhouette Projector first computes the object silhouette heights $H^{\prime}_{\mathrm{obj}}$ from the depth map. By comparing $H^{\prime}_{\mathrm{obj}}$ and the estimated one $H_{\mathrm{obj}}$ with LSP, per-frame scale factor $s$ is determined and we obtain the scaled camera height $H^{n,\tau}_{\mathrm{cam}}=s\cdot H^{\prime n,\tau}_{\mathrm{cam}}$. At the end of each epoch, $H^{*}_{\mathrm{cam}}$ is optimized across a series of consecutive frames and updated with the weighted moving average.
4 FUMET
Fig. 1 depicts the overall architecture of FUMET. Please see the caption for an overview of the framework. Let “$\,{}^{\prime}\,$” denote “not metrically scaled”.
@@ 4.1 Scale-Aware Self-Supervised Learning
In FUMET, we derive a camera height pseudo-supervision from depth estimates. To obtain the per-frame camera height, we first compute the normal vector for each pixel $p_{i}$. We consider the 8-neighborhood of $p_{i}$ and form 8 pairs $\mathcal{N}(p_{i})=\{(p^{i}_{j_{0}},p^{i}_{j_{1}})\}^{8}_{j=1}$. Each pair $(p^{i}_{j_{0}},p^{i}_{j_{1}})$ consists of two pixels whose difference vectors w.r.t. $p_{i}$ are perpendicular to each other and are sorted counterclockwise. We reproject the paired pixels onto 3D space and compute the cross-product of the two difference vectors w.r.t. the reprojected 3D point of $p_{i}$. By summing and normalizing all the cross-product vectors of each pair, we obtain the normal vector $\hat{\mathbf{n}}(p_{i})$
@@
where $p_{i}$ is in the road region mask $M_{\mathrm{r}}$ obtained with an off-the-shelf semantic segmentation model [26] before training. This camera height $\tilde{H}^{\prime}_{\mathrm{cam}}$ estimate is unscaled and needs to be upgraded to a real-scaled value to impose the invariance loss across frames. By using the Silhouette Projector, we aggregate the scale information from the object height prior into a per-frame scale factor $s$ (see Sec. 4.2) and obtain the scaled camera height as $H_{\mathrm{cam}}=s\cdot\tilde{H}^{\prime}_{\mathrm{cam}}$.
We leverage the fact that the camera height does not change in a sequence and formulate the invariance into an optimization across frames and training epochs. Thanks to this optimization, we acquire an accurate and stable supervision, which is also consistent across the scene. At the end of the $\tau$-th epoch, we use the median of $\{H^{n,\tau}_{\mathrm{cam}}\}^{N}_{n=1}$ for all frames $N$ as the representative camera height $H^{\tau}_{\mathrm{cam}}$ at the $\tau$-th epoch. Then, we update the scaled camera height pseudo-invariance $H^{*}_{\mathrm{cam}}$ with its weighted moving average
@@
When training on a dataset containing images captured at multiple camera heights, we create pseudo-labels corresponding to each camera height and optimize them individually.
To make the network learn the absolute scale and directly produce the scaled depth map, we enforce a loss on the camera height $H^{\prime}_{\mathrm{cam}}$ using the pseudo-invariance $H^{*\tau-1}_{\mathrm{cam}}$ computed in the previous epoch
@@
It is worth noting that at inference, the network does not assume camera height invariance in the input sequence; it is only assumed during training. The computational cost at inference is exactly the same as the original MDE model trained without FUMET.